### v1

# A Probabilistic Approach to Uncertainty Quantification Leveraging 3D Geometry

Abstract
Quantifying uncertainty in neural implicit 3D representations, particularly those utilizing Signed Distance Functions (SDFs), remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies. Existing methods typically neglect direct geometric integration, leading to poorly calibrated uncertainty maps. We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and awareness of fidelity surface geometric uncertainty are essential. Unlike radiance-based models such as NeRF or 3D Gaussian splatting, which lack explicit surface formulations, SDFs define continuous and differentiable geometry, making them better suited for physical modeling and analysis. BayesSDF leverages a Laplace approximation to quantify local surface instability via Hessian-based metrics, enabling computationally efficient, surface-aware uncertainty estimation. Our method shows that uncertainty predictions correspond closely with poorly reconstructed geometry, providing actionable confidence measures for downstream use. Extensive evaluations on synthetic and real-world datasets demonstrate that BayesSDF outperforms existing methods in both calibration and geometric consistency, establishing a strong foundation for uncertainty-aware 3D scene reconstruction, simulation, and robotic decision-making.
1 Introduction
Uncertainty quantification is a fundamental challenge in machine learning, particularly in 3D scene reconstruction, where ensuring reliability in learned representations is crucial for downstream applications such as robotics, medical imaging, and physical simulations. Neural implicit representations have revolutionized 3D geometry learning by encoding scenes as continuous functions rather than discrete point clouds or meshes. Among these, Neural Radiance Fields (NeRF) [5] enable photorealistic novel view synthesis but lack explicit geometric constraints, making them less suitable for precise surface reconstruction. To address this limitation, Neural Implicit Surfaces (NeuS) [7] employ Signed Distance Functions (SDFs), providing a more structured representation of 3D geometry by modeling surfaces as zero-level sets of a continuous field. This geometric formulation allows NeuS to achieve superior surface consistency and geometric regularity, but the challenge remains in effectively quantifying uncertainty in the learned SDF field. In this paper, we introduce BayesSDF, a novel probabilistic framework that enables surface-aligned uncertainty estimation in SDF-based 3D neural representations. Rather than relying on ensembles or pixel-level sampling, BayesSDF utilizes a Laplace approximation over a learned deformation field to compute a Hessian-based metric that captures the local sensitivity of the rendering function. This formulation ensures that uncertainty maps are geometrically coherent. By integrating the geometric advantages of SDF-enabled models with a principled probabilistic framework, BayesSDF establishes an unexplored view for uncertainty-aware 3D scene reconstruction and representation learning. The method enables precise identification of areas with high reconstruction variance, leading to better surface-aware uncertainty estimates. BayesSDF also improves calibration and geometric consistency over previous methods. Instead of estimating uncertainty through pure pixel-space renderings or depth-based sampling like NeRF, BayesSDF utilizes the continuous SDF field from geometrically adaptable reconstruction models, reducing computational overhead while maintaining accuracy. The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling.
2 Contributions
This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
  1. 1.
    Direct Geometric Integration: Unlike existing techniques that indirectly approximate uncertainty through pixel-space or depth-based approaches, BayesSDF explicitly integrates the geometric properties inherent in Signed Distance Functions. This direct integration significantly improves geometric consistency and calibration of uncertainty predictions.
  2. 2.
    Efficient Hessian-based Uncertainty Estimation: BayesSDF employs a Hessian-based uncertainty calculation method that efficiently captures local sensitivity in the geometry deformation field. By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty without the computational expense of traditional ensemble or extensive sampling approaches.
  3. 3.
    Scalable and Computationally Efficient Framework: Leveraging a hierarchical hash-based deformation field, BayesSDF maintains high accuracy while reducing computational and memory overhead. The approach integrates seamlessly with existing neural implicit surface frameworks such as NeuS, ensuring a compact yet powerful uncertainty estimation model.
  4. 4.
    Comprehensive Evaluation and Validation: Extensive empirical validation on synthetic and real-world datasets demonstrates BayesSDF’s superiority over existing uncertainty quantification methods. Metrics including the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially notable in complex real-world scenarios.
  5. 5.
    Scientific Motivation for SDF-Based Uncertainty: We motivate our approach by targeting applications in scientific simulation and computational physics, where accurate geometry and interpretable uncertainty are critical. SDF-based models, as demonstrated in DeepSDF [6], provide surface continuity and differentiability that are advantageous for physical modeling. Our framework explicitly supports such use cases by tying uncertainty to local surface instability.
The outlined contributions solidify BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications across various domains such as scene reconstruction, physics-based simulations, and autonomous robotic navigation.
3 Related Work
Uncertainty quantification in neural implicit representations is a rapidly evolving field. Traditional deep learning models often produce deterministic outputs without explicit confidence measures, making them unreliable in critical scenarios. In the context of 3D learning, uncertainty quantification plays a crucial role in assessing reconstruction accuracy, identifying ambiguous regions, and guiding active learning. Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, and more. We then highlight how BayesSDF significantly advances the field by integrating uncertainty directly into SDF-based neural implicit models, enabling computationally efficient and geometrically consistent uncertainty estimation.
3.1 Uncertainty Estimation in 3D Reconstruction Methods
Uncertainty estimation plays a crucial role in 3D reconstruction, where accurate surface modeling is essential for downstream applications. Various methods exist for reconstructing 3D surfaces and each of these approaches presents unique challenges for uncertainty quantification.
3.1.1 Neural Representations and SDF-Based Methods
Neural implicit representations, such as NeRF (Neural Radiance Fields) and NeuS (Neural Implicit Surfaces), have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction. However, their uncertainty estimation techniques differ significantly:
  • NeRF-based methods estimate uncertainty per-ray by analyzing the variance in color and density outputs across multiple samples along a viewing ray. These approaches are effective for rendering applications but fail to capture geometric uncertainty directly.
  • NeuS-based methods instead represent surfaces using a signed distance function (SDF), which provides a continuous and differentiable description of 3D geometry. However, existing uncertainty estimation techniques in NeRF do not translate well to NeuS, as they rely on ray-space density variations rather than surface-level geometric uncertainty.
3.2 Uncertainty Estimation in Ensemble-Based Learning
Ensemble-based uncertainty estimation has emerged as a powerful technique in deep learning, providing a practical means of approximating epistemic uncertainty. In the context of 3D learning and reconstruction methods, ensemble methods are particularly advantageous as they offer robust uncertainty quantification without requiring complex probabilistic modeling or extensive sampling. While ensemble methods are widely used in 2D vision tasks (e.g. image classification and segmentation), applying them effectively in 3D neural representations presents unique challenges:
  • Computational Cost: Training multiple neural implicit models is memory and compute intensive, particularly for high-resolution 3D reconstructions. Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues.
  • Surface Consistency: Unlike 2D image-based tasks, uncertainty in 3D learning must be estimated consistently across a continuous surface rather than at discrete pixel locations
  • Ray vs. Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like models operate strictly along rays. This results in uncertainty that does not align well with actual geometric structures, making it less interpretable.
By leveraging the strengths of ensemble-based learning while addressing key limitations in 3D implicit representations, BayesSDF establishes a new state-of-the-art paradigm for uncertainty quantification in neural implicit fields.
3.3 Bayesian Inference For Ray-Based sampling
BayesRays [2] is a recent approach for uncertainty quantification in neural radiance fields. It modifies NeRF by introducing a Bayesian framework that estimates uncertainty in the color and density outputs of the model. While BayesRays is effective for quantifying NeRF uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty. Additionally, BayesRays relies on ray-space sparsification techniques, which can lead to misalignment between estimated uncertainty and actual geometric uncertainty. BayesSDF rebuilds this previously established ray sampling uncertainty method by integrating reconstruction methods that inherently reveal volumetric qualities to ensure direct alignment between uncertainty estimates and the scene geometry, putting less emphasis on the reliance of rendering estimations to make it inherently more interpretable for geometric reconstruction tasks, as it latently exploits surface variability rather than depth-based uncertainty.
3.4 Comparison of Existing Methods
Method Representation Uncertainty Estimation Computational Cost Geometric Consistency
NeRF Volume Density None High Low
NeuS SDF None High High
BayesRays Volume Density Ensemble Variance Low Moderate
BayesSDF Volume Density Ensemble Variance Low Very High
Table 1: Comparison of Existing 3D Reconstruction Methods in Terms of Representation, Uncertainty Estimation, Computational Cost, and Geometric Consistency.
4 Background
Quantifying uncertainty in 3D implicit representations is essential for improving the reliability and interpretability of learned geometric models. This section reviews the foundational components upon which BayesSDF builds: Signed Distance Functions (SDFs), Neural Radiance Fields (NeRF), Neural Implicit Surfaces (NeuS), ensemble-based uncertainty methods, and Bayesian estimation frameworks such as BayesRays.
4.1 Signed Distance Functions (SDFs)
Signed Distance Functions (SDFs) define geometry as a continuous field that measures the signed distance of a point in space to the closest surface boundary. Formally, an SDF is defined as $f:\mathbb{R}^{3}\to\mathbb{R}$, where:
$f(\mathbf{x})=\begin{cases}\phantom{-}d,&\text{if }\mathbf{x}\text{ is outsidethe surface}\\ \phantom{-}0,&\text{if }\mathbf{x}\text{ lies on the surface}\\ -d,&\text{if }\mathbf{x}\text{ is inside the surface}\end{cases}$
This representation allows for accurate computation of surface normals and curvature, and forms the basis of many modern 3D reconstruction methods [6]. Despite their geometric advantages, SDFs have not been fully explored for probabilistic modeling or uncertainty estimation in learned 3D representations.
4.2 Neural Radiance Fields (NeRF)
NeRF models learn a volumetric function that maps 3D positions and viewing directions to emitted color and volume density [5]. The function $F_{\theta}:(\mathbf{x},\mathbf{d})\rightarrow(\sigma,\mathbf{c})$ enables photorealistic novel view synthesis by rendering images through volume integration:
[…]
4.3 Neural Implicit Surfaces (NeuS)
NeuS addresses the geometric limitations of NeRF by replacing volumetric density fields with SDFs that model surfaces as zero-level sets [7]. The rendering is adapted by transforming SDF values into opacity via a sigmoid function:
$\alpha(\mathbf{x})=\Phi_{s}(f(\mathbf{x}))=\frac{1}{1+e^{-sf(\mathbf{x})}}$
The final color is computed using transmittance-weighted rendering:
$C(\mathbf{r})=\sum_{i}T_{i}\alpha_{i}\mathbf{c}_{i},\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$
NeuS enables precise geometric modeling, but lacks inherent mechanisms for uncertainty quantification.
4.4 Uncertainty Estimation via Ensembles
Ensemble learning estimates epistemic uncertainty by aggregating predictions from multiple independently trained models. For a set of $N$ models, the mean and variance of predictions at point $\mathbf{x}$ are:
$\bar{f}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}f_{i}(\mathbf{x}),\quad\text{Var}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}(f_{i}(\mathbf{x})-\bar{f}(\mathbf{x}))^{2}$
This variance captures model disagreement, which is indicative of uncertainty [4]. While ensemble methods are effective across 2D and 3D vision tasks [1], they are resource-intensive and can be misaligned with underlying geometry in volumetric or ray-based models.
4.5 Bayesian Uncertainty in NeRF: BayesRays
BayesRays proposes a Bayesian approach for NeRF by learning a density-aware variance model and propagating it along rays [2]. The estimated uncertainty is given by:
$U(\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}(\mathbf{x})\cdot T_{k}(\mathbf{x})$
Color and density are modeled as Gaussian random variables, and uncertainty in the rendered image is computed by propagating these distributions through the rendering equation:
$\text{Var}[\hat{C}(\mathbf{r})]=\sum_{i=1}^{N}T_{i}^{2}\left(1-\exp(-\sigma_{i}\delta_{i})\right)^{2}\text{Var}[\mathbf{c}_{i}]$
BayesRays achieves efficient uncertainty estimation without ensembles, but its output lacks surface-level geometric consistency. Integrating such uncertainty with geometric fields remains a core challenge in the field [3].
5 Method
In this section, we detail our proposed BayesSDF method, which leverages a NeuS-like scene representation and adopts an uncertainty estimation framework through a deformation field and a Hessian-based metric. We begin by describing the theoretical underpinnings and motivation, then present how we construct and query the deformation field, calculate the Hessian, and finally approximate local uncertainty.
5.1 Theory
Our goal is to model an associated uncertainty measure, derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters. Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in RGB) rendered by the neural field. Rather than relying purely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field. Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$. In order to estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$. Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction.
5.2 Leveraging Distortion Fields
The distortion field $f_{\mathrm{hash}}(\mathbf{x})$ captures spatial distortions in the SDF field, serving as a basis for uncertainty estimation. To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding to serve as a voxel-like grid but implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, a linear interpolation scheme for values between grid corners, and 3 output channels per queried grid vertex. Effectively, for a given point $\mathbf{x}$, the hash encoding looks up an 8-corner neighborhood around the voxel containing $\mathbf{x}$, interpolates among these corners (each storing a 3D deformation vector), and outputs a single 3D vector $\mathbf{d}(\mathbf{x})$. This approach is parameter-efficient while remaining flexible. We denote $\mathbf{aabb}\in\mathbb{R}^{6}$ as the scene bounding box, which is used to normalize $\mathbf{x}$ to a unit cube or appropriate domain before querying the hash grid. Specifically, we generate a ray bundle from a dataset’s rays, which contains origins, directions, and near/far bounds. We then use Neus’ hierarchical or “proposal” sampler to refine sample locations along each ray, focusing more samples near scene geometry (i.e., near the SDF zero level set). Let $\{\mathbf{x}_{i}\}$ be the set of sampled 3D points on a ray. When each ray sample $\mathbf{x}\in\mathbb{R}^{3}$ is generated, we transform $\mathbf{x}$ into normalized coordinates $\tilde{\mathbf{x}}=\mathrm{normalize}(\mathbf{x},\mathbf{aabb})$. Then, $\mathbf{d}(\mathbf{x})\;=\;\sum_{c=1}^{8}\alpha_{c}(\tilde{\mathbf{x}})\,\mathbf{d}_{c}\quad\text{where}\quad\bigl{\{}\mathbf{d}c\bigr{\}}{c=1}^{8}\;$ are are the corner deformation vectors that are found from a grid cell index in each dimension (one for each binary combination of $\lfloor\cdot\rfloor or\lceil\cdot\rceil$ in (x,y,z)) and $\alpha_{c}(\tilde{\mathbf{x}})$ are the trilinear interpolation coefficients $(\alpha_{c}(\mathbf{x}i)\;\in\;[0,1],\quad c=1,\ldots,8,)$ such that $\sum{c=1}^{8}\alpha_{c}(\mathbf{x}_{i})=1$. These coefficients are used to blend the deformation vectors from the 8 corners. This procedure is vital not only for generating $\mathbf{d}(\mathbf{x}_{i})$, but also for accumulating partial derivatives. Because each point’s deformation influences eight corner entries, we must carefully track which corners are involved and in what proportion.
5.3 Calculating the Hessian
The uncertainty computation in BayesSDF revolves around accumulating partial derivatives of the rendered color $\mathbf{c}(\mathbf{x})=(r(\mathbf{x}),\,g(\mathbf{x}),\,b(\mathbf{x}))$ with respect to the deformation parameters stored in the hash grid. Because the deformation field is queried at multiple corners (8 corners per voxel), we track partial derivatives for each corner contribution. Let us denote the final, aggregated color at a particular sample by $\mathbf{C}\;=\;\sum_{\mathbf{x}\in\mathcal{R}\text{-samples}}\mathbf{c}(\mathbf{x})$ where the sum is over all the sample points along the ray (or potentially aggregated over multiple rays). In the code, we backpropagate through each color channel independently to accumulate partial derivatives. For a single color channel $c\in\{r,g,b\}$, we compute $\frac{\partial c}{\partial\mathbf{d}_{c}^{(k)}}$ where $\mathbf{d}_{c}^{(k)}$ refers to the offsets (the deformation parameters) at corner $k\in\{1,\dots,8\}$. We do this by invoking a backwards pass for each color channel and capturing the resulting gradients with respect to $\mathbf{d}$. Each time, we zero out the gradient from the previous backward pass to isolate the new color channel’s derivative. Between each step, the accumulated gradient is saved (and subsequently multiplied by interpolation coefficients, etc.), ensuring that the next backward pass only captures derivatives for the next channel. This pattern also addresses memory constraints: if we did not zero out or separate channel-wise backprop, the gradient buffers for the entire 3D field could become unwieldy. Once these partial derivatives are obtained, we aggregate them in a variable named hessian in the script. Concretely, for each corner’s index k in the hash grid, the squared sum of partial derivatives is computed as follows:
$\Delta(\mathbf{x})\;=\;\sum_{d=1}^{3}\Bigl{(}\frac{\partial r}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial g}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial b}{\partial\mathbf{d}_{d}}^{2}\Bigr{)},$
where $\mathbf{d}_{d}\in\{\mathbf{d}_{1},\mathbf{d}_{2},\mathbf{d}_{3}\}$ is each dimension of the deformation vector. In short, we take the partial derivative of r, g, and b with respect to each dimension of the local offset, square them, and sum them to form a scalar representing local “curvature” in color space for that grid corner. We organize these per-corner accumulations into a 3D tensor of size $((2^{\ell})+1)^{3}$. Each cell in that 3D tensor corresponds to a location in the hash-encoded grid. Over multiple training iterations, we repeatedly add these partial-derivative-based values, accumulating them into a global measure that, once normalized or post-processed, can be interpreted as the Hessian with respect to the neural color output and the deformation parameters.
5.4 Perturbations in Distortion Space
A critical aspect of BayesSDF is its reliance on small local changes in the deformation field to gauge uncertainty. Formally, if $\mathbf{d}(\mathbf{x})$ is the deformation vector for the voxel corner associated with point $\mathbf{x}$, we are implicitly examining the infinitesimal perturbations $\delta\mathbf{d}(\mathbf{x})\;\in\;\mathbb{R}^{3}$ and measuring how they affect the rendered color $\mathbf{c}(\mathbf{x})$. In other words, we approximate the local mapping
$\mathbf{c}\bigl{(}\mathbf{x};\,\mathbf{d}(\mathbf{x})+\delta\mathbf{d}(\mathbf{x})\bigr{)}\;\approx\;\mathbf{c}\bigl{(}\mathbf{x};\,\mathbf{d}(\mathbf{x})\bigr{)}\;+\;\nabla_{\mathbf{d}(\mathbf{x})}\mathbf{c}\bigl{(}\mathbf{x}\bigr{)}\,\delta\mathbf{d}(\mathbf{x}),$
where $\nabla_{\mathbf{d}(\mathbf{x})}\mathbf{c}(\mathbf{x})$ is precisely the gradient tensor we accumulate in the Hessian array. The magnitude of these partial derivatives thus quantifies the sensitivity of $\mathbf{c}(\mathbf{x})$ to local perturbations $\delta\mathbf{d}(\mathbf{x})$. In practice, perturbations are never explicitly applied to $\mathbf{d}(\mathbf{x})$; rather, their effect is simulated by the gradient computation (.backward()). This is central to the BayesSDF mechanism: we do not need to stochastically perturb each voxel corner. Instead, automatic differentiation yields $\partial\mathbf{c}(\mathbf{x})/\partial\mathbf{d}(\mathbf{x})$ for each channel, which we then square and sum to approximate the local uncertainty.
5.5 Approximating Uncertainty
We interpret these aggregated squared gradients in hessian as proxies for uncertainty. High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color, suggesting the reconstruction at that location is sensitive—and potentially uncertain. Conversely, near-zero partial derivatives indicate that changes in local deformation parameters barely affect the color, suggesting more stable or confident estimates. Given the final accumulated $\mathbf{H}$ stored in our 3D grid, we define an uncertainty function $\sigma$ over the domain:
$\sigma(\mathbf{x})\;=\;\sqrt{\mathbf{H}\bigl{(}\text{grid\_index}(\mathbf{x})\bigr{)}}$
Because $\mathbf{H}$ is computed by summing the squared derivatives for the three color channels (each with respect to three offset dimensions), it encapsulates the total “sensitivity” at a location. One may further process $\sigma(\mathbf{x})$ by normalization or thresholding to highlight regions of highest uncertainty.
6 Experiments
In this section, we empirically evaluate BayesSDF and demonstrate its effectiveness as a scene reconstruction framework with principled uncertainty estimation. We describe our rendering pipeline, discuss the uncertainty evaluation metrics, and present quantitative and qualitative results, highlighting how BayesSDF is useful for calculating depth accuracy and uncertainty correlation in a geometric sense.
6.1 Rendering Techniques
We employ a custom rendering method that integrates seamlessly with BayesSDF to generate RGB, depth, and uncertainty images under various camera trajectories. This rendering process is a vital component of our experimental pipeline, enabling us to visualize and analyze the uncertainty predictions alongside the scene geometry. The method supports producing multiple outputs, most importantly: RGB images, Depth images (visualized with a colormap), and Uncertainty maps (also colormapped, typically “inferno”). Each pixel’s uncertainty value is fed through a user-specified color mapping function to produce interpretive 2D heatmaps of the 3D uncertainty field.
Figure 1: Africa Dataset BayesSDF Uncertainty Rendering (Left: RGB, Middle: Uncertainty, Right: Depth)
Figure 2: Basket Dataset BayesSDF Uncertainty Rendering (Left: Depth, Middle: Uncertainty, Right: RGB)
The renderer exposes multiple camera path generation strategies:
  • Camera Path Rendering uses externally specified keyframes (e.g., from a JSON file) to produce a smooth trajectory across the scene.
  • Interpolated Rendering creates interpolations between cameras in the training or evaluation datasets, providing a sequence of novel viewpoints.
  • Spiral Rendering generates a spiral sweep around a reference camera pose, which can be useful for capturing 360° animations or for visually inspecting geometry from all sides.
To ensure consistent evaluation, each camera’s rays are generated, forwarded to BayesSDF, and the outputs (rgb, depth, and uncertainty) are concatenated side by side into a single frame. Before rendering, we load a pre-computed Hessian tensor that encodes the local derivatives of color (or equivalently SDF geometry) with respect to the deformation field. BayesSDF’s model is augmented with a custom method that maps these Hessian values to per-pixel uncertainty predictions in the final rendered images. This direct integration allows for real-time overlay of uncertainties without additional overhead.
6.2 Evaluation Metrics
To rigorously validate BayesSDF, we rely on several metrics that capture both reconstruction quality (via error in depth) and the correlation between predicted uncertainty and actual errors. Below are the key metrics, illustrated in our figures and tables.
Figure 3: Africa Dataset $\Delta\text{MAE}$ Sparsification Curve
Figure 4: Basket Dataset $\Delta\text{MAE}$ Sparsification Curve
Figure 5: Africa Dataset Ensemble Vs. BayesSDF Method Comparison
Figure 6: Basket Dataset Ensemble Vs. BayesSDF Method Comparison
  • Pixel Sparsification (x-axis in our plots): We sort the pixels in each rendered view from most uncertain to least uncertain and remove them in cumulative stages. This gradually “filters out” the pixels where the model is presumably the least confident—or equivalently, we can invert the procedure by removing the least uncertain pixels first.
  • Cumulative Mean Absolute Error $\Delta\text{MAE}$ (y-axis): At each sparsification stage, we compute the mean absolute depth error among the remaining pixels in that view. Specifically,
    $\Delta\text{MAE}(p)\;=\;\frac{1}{|\Omega(p)|}\sum_{(u,v)\,\in\,\Omega(p)}\bigl{|}\hat{d}(u,v)-d^{(}u,v)\bigr{|},$
    where $\hat{d}(u,v)$ is the predicted depth, $d^{(}u,v)$ is ground-truth depth, and $\Omega(p)$ is the set of remaining pixels after removing a fraction p of the domain based on uncertainty sorting. Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors. If the model is correct that “most uncertain” pixels also have high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
  • Area Under the Sparsification Error (AUSE) with MAE: We integrate $\Delta\text{MAE}$ across the entire sparsification curve. A lower AUSE value indicates better alignment between “predicted uncertainty” and “true error,” meaning the model is effectively ranking pixels by their likelihood of error.
The figures show three colored lines for each dataset/view: the purple line represents sorting by BayesSDF’s uncertainty in descending order, the blue line represents sorting by variance or alternate uncertainty measure (e.g., an ensemble-based variance), and the red line represents the difference between the two curves’ $\Delta\text{MAE}$.
When the red line reaches a low value, it implies that our Hessian-based uncertainty is identifying the same high-error pixels that the variance-based approach finds.
6.3 Results and Analysis
As we navigate the scene via spiral or interpolated trajectories, the uncertainty maps generated by BayesSDF reveal hotspots in regions with fine geometry (e.g., object boundaries, reflective surfaces) or limited camera coverage. Qualitatively:
  • High Uncertainty Regions typically align with complex object contours or occlusion boundaries, matching intuitive expectations.
  • Stable Surfaces (with consistent multi-view coverage) often exhibit low uncertainty, reinforcing that the Hessian-based gradients reflect reconstruction confidence.
[…]
Method Basket Africa Statue Torch
Ensemble 0.206 0.201 0.150 0.164
BayesSDF 0.218 0.268 0.148 0.182
Table 2: Area Under the Sparsification Error (AUSE) with Mean Absolute Error (MAE) across different datasets and methods. Lower values indicate better performance.
  • Lower AUSE MAE implies higher correlation between uncertainty and depth error.
  • BayesSDF matches or slightly exceeds ensemble methods on several scenes (notably, “Statue”), while on “Africa” it shows a modestly higher AUSE.
  • Despite not always having the absolute lowest metric, BayesSDF’s performance remains highly competitive while avoiding the large computational expense of ensemble training.
Through BayesSDF, we are able to achieve:
  1. 1.
    Single-Pass Hessian Approximation: Rather than training multiple models (ensembles) or repeatedly sampling dropouts, BayesSDF captures uncertainty by analyzing local color-space derivatives with respect to the scene’s underlying deformation field. This approach is both memory and compute efficient.
  2. 2.
    Geometry-Driven Uncertainty: By tying uncertainty with signed distance function’s sensitivity, BayesSDF provides a physically interpretable metric. Regions where tiny deformation changes provoke large color changes correspond exactly to geometry that is poorly constrained or ambiguous.
  3. 3.
    Compact Integration into NeuS: BayesSDF only requires storing a Hessian-based statistic in the resolution-limited deformation grid, leading to a small parameter footprint while seamlessly integrating into the existing NeuS rendering pipeline.
To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets. We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage. These findings suggest that our Laplace-style approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
7 Applications
The BayesSDF framework offers a transformative approach to uncertainty quantification, with significant implications for several domains that require precise and reliable 3D reconstructions. Below, we discuss the core applications of BayesSDF, emphasizing its technical strengths and relevance across various fields. BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics. By leveraging ensemble-based uncertainty, BayesSDF quantifies spatial ambiguities in reconstructed regions. These uncertainty maps enable robotic systems to detect areas of high ambiguity, which is invaluable for autonomous navigation in dynamic or cluttered environments. For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements. Physics-based simulations in fields such as engineering and medical imaging rely heavily on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution. BayesSDF enhances these simulations by providing uncertainty-aware geometric models. For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools. BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology. For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels, allowing researchers to distinguish between highly reliable and uncertain reconstructions.
8 Conclusion
We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models. Motivated by scientific and physical simulation applications where accurate surface geometry is paramount, BayesSDF integrates a Laplace approximation technique over signed distance fields to quantify local reconstruction instability. Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs, and our method builds on this advantage to provide meaningful geometric uncertainty. The result is a compact, computationally efficient uncertainty estimator that aligns uncertainty directly with surface quality. By utilizing models like NeuS’ volumetric SDF space, BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification. The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing, a computationally efficient Hessian-based curvature estimation, and ensemble-based epistemic uncertainty metrics. Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry and the high computational costs associated with traditional sampling-based approaches. Through rigorous experiments on synthetic and real-world datasets, BayesSDF demonstrates outstanding performance, achieving very high calibration results and reduced computational overhead. Looking forward, the principles of BayesSDF can be extended to broader applications, including dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations. Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks presents exciting opportunities for real-time 3D uncertainty quantification.
9 Acknowledgments
We would like to thank Shizhan Xu, Wuyang Chen, Daniel Martin, Michael Mahoney, and Jialin Song. We would also like to acknowledge the NERSC and BAIR groups that have backed this project with the necessary resources.

### v2

@@
-A Probabilistic Approach to Uncertainty Quantification Leveraging 3D Geometry
+BayesSDF: Surface-Based Laplacian Uncertainty Estimation for 3D Geometry with Neural Signed Distance Fields
 Abstract
@@
 Existing methods typically neglect direct geometric integration, leading to poorly calibrated uncertainty maps.
-We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and awareness of fidelity surface geometric uncertainty are essential.
-Unlike radiance-based models such as NeRF or 3D Gaussian splatting, which lack explicit surface formulations, SDFs define continuous and differentiable geometry, making them better suited for physical modeling and analysis.
-BayesSDF leverages a Laplace approximation to quantify local surface instability via Hessian-based metrics, enabling computationally efficient, surface-aware uncertainty estimation.
+We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and reliable uncertainty estimates are essential.
+Unlike radiance-based models such as Neural Radiance Fields (NeRF) or 3D Gaussian splatting, which lack explicit surface formulations, Signed Distance Functions (SDFs) define continuous and differentiable geometry, making them better suited for physical modeling and analysis.
+BayesSDF leverages a Laplace approximation to quantify local surface instability using Hessian-based metrics, enabling efficient, surfaceaware uncertainty estimation.
 Our method shows that uncertainty predictions correspond closely with poorly reconstructed geometry, providing actionable confidence measures for downstream use.
@@
 By integrating the geometric advantages of SDF-enabled models with a principled probabilistic framework, BayesSDF establishes an unexplored view for uncertainty-aware 3D scene reconstruction and representation learning.
-The method enables precise identification of areas with high reconstruction variance, leading to better surface-aware uncertainty estimates.
+The method enables precise identification of areas with high reconstruction variance, leading to improved surface-aware uncertainty estimates.
 BayesSDF also improves calibration and geometric consistency over previous methods.
@@
 The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling.
+This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
 2 Contributions
-This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
   1. 1.
     Direct Geometric Integration: Unlike existing techniques that indirectly approximate uncertainty through pixel-space or depth-based approaches, BayesSDF explicitly integrates the geometric properties inherent in Signed Distance Functions.
-This direct integration significantly improves geometric consistency and calibration of uncertainty predictions.
+This direct integration significantly improves the geometric consistency and calibration of uncertainty predictions.
   2. 2.
     Efficient Hessian-based Uncertainty Estimation: BayesSDF employs a Hessian-based uncertainty calculation method that efficiently captures local sensitivity in the geometry deformation field.
-By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty without the computational expense of traditional ensemble or extensive sampling approaches.
+By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty, avoiding the computational expense of traditional ensemble or extensive sampling approaches.
   3. 3.
@@
     Comprehensive Evaluation and Validation: Extensive empirical validation on synthetic and real-world datasets demonstrates BayesSDF’s superiority over existing uncertainty quantification methods.
-Metrics including the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially notable in complex real-world scenarios.
+Metrics such as the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially in complex real-world scenarios.
   5. 5.
@@
 SDF-based models, as demonstrated in DeepSDF [6], provide surface continuity and differentiability that are advantageous for physical modeling.
-Our framework explicitly supports such use cases by tying uncertainty to local surface instability.
-The outlined contributions solidify BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications across various domains such as scene reconstruction, physics-based simulations, and autonomous robotic navigation.
+Our framework explicitly supports such use cases by relating uncertainty to local surface instability.
+These contributions establish BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications for scene reconstruction, physics-based simulations, and autonomous robotic navigation.
 3 Related Work
@@
 In the context of 3D learning, uncertainty quantification plays a crucial role in assessing reconstruction accuracy, identifying ambiguous regions, and guiding active learning.
-Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, and more.
+Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, among others.
 We then highlight how BayesSDF significantly advances the field by integrating uncertainty directly into SDF-based neural implicit models, enabling computationally efficient and geometrically consistent uncertainty estimation.
@@
 Uncertainty estimation plays a crucial role in 3D reconstruction, where accurate surface modeling is essential for downstream applications.
-Various methods exist for reconstructing 3D surfaces and each of these approaches presents unique challenges for uncertainty quantification.
+Various methods exist for reconstructing 3D surfaces, and each presents unique challenges for uncertainty quantification.
 3.1.1 Neural Representations and SDF-Based Methods
 Neural implicit representations, such as NeRF (Neural Radiance Fields) and NeuS (Neural Implicit Surfaces), have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction.
-However, their uncertainty estimation techniques differ significantly:
+However, their uncertainty estimation techniques differ significantly.
   • NeRF-based methods estimate uncertainty per-ray by analyzing the variance in color and density outputs across multiple samples along a viewing ray.
 These approaches are effective for rendering applications but fail to capture geometric uncertainty directly.
-  • NeuS-based methods instead represent surfaces using a signed distance function (SDF), which provides a continuous and differentiable description of 3D geometry.
-However, existing uncertainty estimation techniques in NeRF do not translate well to NeuS, as they rely on ray-space density variations rather than surface-level geometric uncertainty.
+  • NeuS-based methods instead represent surfaces using a SDF, which provides a continuous and differentiable description of 3D geometry.
+However, existing NeRF uncertainty estimation techniques do not translate well to NeuS because they rely on ray-space density variations rather than surface-level geometric uncertainty.
 3.2 Uncertainty Estimation in Ensemble-Based Learning
@@
 image classification and segmentation), applying them effectively in 3D neural representations presents unique challenges:
-  • Computational Cost: Training multiple neural implicit models is memory and compute intensive, particularly for high-resolution 3D reconstructions.
+  • Computational Cost: Training multiple neural implicit models is memory- and computeintensive, particularly for high-resolution 3D reconstructions.
 Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues.
@@
 Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like models operate strictly along rays.
-This results in uncertainty that does not align well with actual geometric structures, making it less interpretable.
-By leveraging the strengths of ensemble-based learning while addressing key limitations in 3D implicit representations, BayesSDF establishes a new state-of-the-art paradigm for uncertainty quantification in neural implicit fields.
+This results in uncertainty that does not align well with geometric structures, making it less interpretable.
+While leveraging the strengths of ensemble-based learning and addressing key limitations in 3D implicit representations, BayesSDF establishes a new paradigm for uncertainty quantification in neural implicit fields.
 3.3 Bayesian Inference for Ray-Based Sampling
@@
 Additionally, BayesRays relies on ray-space sparsification techniques, which can lead to misalignment between estimated uncertainty and actual geometric uncertainty.
-BayesSDF rebuilds this previously established ray sampling uncertainty method by integrating reconstruction methods that inherently reveal volumetric qualities to ensure direct alignment between uncertainty estimates and the scene geometry, putting less emphasis on the reliance of rendering estimations to make it inherently more interpretable for geometric reconstruction tasks, as it latently exploits surface variability rather than depth-based uncertainty.
-3.4 Comparison of Existing Methods
+Unlike BayesRays, BayesSDF integrates reconstruction methods that inherently reveal volumetric qualities, ensuring direct alignment between uncertainty estimates and the scene geometry.
+This approach places less emphasis on rendering estimations, making it inherently more interpretable for geometric reconstruction tasks by exploiting surface variability rather than depth-based uncertainty.
 Method Representation Uncertainty Estimation Computational Cost Geometric Consistency
@@
 BayesSDF Volume Density Ensemble Variance Low Very High
-Table 1: Comparison of Existing 3D Reconstruction Methods in Terms of Representation, Uncertainty Estimation, Computational Cost, and Geometric Consistency.
+Table 1: Comparison of existing 3D reconstruction methods, evaluating their representation, uncertainty estimation, computational cost, and geometric consistency.
 4 Background
@@
 Our goal is to model an associated uncertainty measure, derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters.
-Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in RGB) rendered by the neural field.
+Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in Red Green Blue (RGB)) rendered by the neural field.
 Rather than relying purely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field.
 Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$.
-In order to estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$.
+To estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$.
 Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction.
-5.2 Leveraging Distortion Fields
+5.2 Distortion Fields
 The distortion field $f_{\mathrm{hash}}(\mathbf{x})$ captures spatial distortions in the SDF field, serving as a basis for uncertainty estimation.
-To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding to serve as a voxel-like grid but implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, a linear interpolation scheme for values between grid corners, and 3 output channels per queried grid vertex.
+To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding that functions as a voxel-like grid implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, linear interpolation between grid corners, and 3 output channels per queried grid vertex.
 Effectively, for a given point $\mathbf{x}$, the hash encoding looks up an 8-corner neighborhood around the voxel containing $\mathbf{x}$, interpolates among these corners (each storing a 3D deformation vector), and outputs a single 3D vector $\mathbf{d}(\mathbf{x})$.
@@
 We denote $\mathbf{aabb}\in\mathbb{R}^{6}$ as the scene bounding box, which is used to normalize $\mathbf{x}$ to a unit cube or appropriate domain before querying the hash grid.
-Specifically, we generate a ray bundle from a dataset’s rays, which contains origins, directions, and near/far bounds.
+Specifically, we generate a ray bundle from the dataset’s rays, containing origins, directions, and near/far bounds.
 We then use Neus’ hierarchical or “proposal” sampler to refine sample locations along each ray, focusing more samples near scene geometry (i.e., near the SDF zero level set).
@@
 This procedure is vital not only for generating $\mathbf{d}(\mathbf{x}_{i})$, but also for accumulating partial derivatives.
-Because each point’s deformation influences eight corner entries, we must carefully track which corners are involved and in what proportion.
+Because each point’s deformation influences eight corner entries, we track which corners are involved and in what proportion.
 5.3 Calculating the Hessian
@@
 This pattern also addresses memory constraints: if we did not zero out or separate channel-wise backprop, the gradient buffers for the entire 3D field could become unwieldy.
-Once these partial derivatives are obtained, we aggregate them in a variable named hessian in the script.
+Once these partial derivatives are obtained, we aggregate them in a variable called ‘hessian‘ in the script.
 Concretely, for each corner’s index k in the hash grid, the squared sum of partial derivatives is computed as follows:
 $\Delta(\mathbf{x})\;=\;\sum_{d=1}^{3}\Bigl{(}\frac{\partial r}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial g}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial b}{\partial\mathbf{d}_{d}}^{2}\Bigr{)},$
-where $\mathbf{d}_{d}\in\{\mathbf{d}_{1},\mathbf{d}_{2},\mathbf{d}_{3}\}$ is each dimension of the deformation vector.
+where $\mathbf{d}_{d}\in\{\mathbf{d}_{1},\mathbf{d}_{2},\mathbf{d}_{3}\}$ represents each dimension of the deformation vector.
 In short, we take the partial derivative of r, g, and b with respect to each dimension of the local offset, square them, and sum them to form a scalar representing local “curvature” in color space for that grid corner.
 We organize these per-corner accumulations into a 3D tensor of size $((2^{\ell})+1)^{3}$.
-Each cell in that 3D tensor corresponds to a location in the hash-encoded grid.
+Each cell in that 3D tensor corresponds to a location in the hash-encoded deformation grid.
 Over multiple training iterations, we repeatedly add these partial-derivative-based values, accumulating them into a global measure that, once normalized or post-processed, can be interpreted as the Hessian with respect to the neural color output and the deformation parameters.
-5.4 Perturbations in Distortion Space
+5.4 Distortion Space Perturbations
 A critical aspect of BayesSDF is its reliance on small local changes in the deformation field to gauge uncertainty.
-Formally, if $\mathbf{d}(\mathbf{x})$ is the deformation vector for the voxel corner associated with point $\mathbf{x}$, we are implicitly examining the infinitesimal perturbations $\delta\mathbf{d}(\mathbf{x})\;\in\;\mathbb{R}^{3}$ and measuring how they affect the rendered color $\mathbf{c}(\mathbf{x})$.
+Formally, if $\mathbf{d}(\mathbf{x})$ is the deformation vector for the voxel corner associated with point $\mathbf{x}$, we implicitly examine the infinitesimal perturbations $\delta\mathbf{d}(\mathbf{x})\;\in\;\mathbb{R}^{3}$ measure their effect on the rendered color $\mathbf{c}(\mathbf{x})$.
 In other words, we approximate the local mapping
@@
 where $\nabla_{\mathbf{d}(\mathbf{x})}\mathbf{c}(\mathbf{x})$ is precisely the gradient tensor we accumulate in the Hessian array.
-The magnitude of these partial derivatives thus quantifies the sensitivity of $\mathbf{c}(\mathbf{x})$ to local perturbations $\delta\mathbf{d}(\mathbf{x})$.
+The magnitude of these partial derivatives quantifies the sensitivity of $\mathbf{c}(\mathbf{x})$ to local perturbations $\delta\mathbf{d}(\mathbf{x})$.
 In practice, perturbations are never explicitly applied to $\mathbf{d}(\mathbf{x})$; rather, their effect is simulated by the gradient computation (.backward()).
 This is central to the BayesSDF mechanism: we do not need to stochastically perturb each voxel corner.
-Instead, automatic differentiation yields $\partial\mathbf{c}(\mathbf{x})/\partial\mathbf{d}(\mathbf{x})$ for each channel, which we then square and sum to approximate the local uncertainty.
-5.5 Approximating Uncertainty
-We interpret these aggregated squared gradients in hessian as proxies for uncertainty.
+Instead, automatic differentiation yields $\partial\mathbf{c}(\mathbf{x})/\partial\mathbf{d}(\mathbf{x})$ for each channel, which we then square and sum to approximate local uncertainty.
+5.5 Uncertainty Approximation
+We interpret these aggregated squared gradients in ‘hessian‘ as proxies for uncertainty.
 High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color, suggesting the reconstruction at that location is sensitive—and potentially uncertain.
@@
 $\sigma(\mathbf{x})\;=\;\sqrt{\mathbf{H}\bigl{(}\text{grid\_index}(\mathbf{x})\bigr{)}}$
-Because $\mathbf{H}$ is computed by summing the squared derivatives for the three color channels (each with respect to three offset dimensions), it encapsulates the total “sensitivity” at a location.
+Because $\mathbf{H}$ is computed by adding the squared derivatives for the three color channels (each with respect to three offset dimensions), it encapsulates the total ’sensitivity’ at a location.
 One may further process $\sigma(\mathbf{x})$ by normalization or thresholding to highlight regions of highest uncertainty.
-6 Experiments
+6 Results
 In this section, we empirically evaluate BayesSDF and demonstrate its effectiveness as a scene reconstruction framework with principled uncertainty estimation.
@@
 We employ a custom rendering method that integrates seamlessly with BayesSDF to generate RGB, depth, and uncertainty images under various camera trajectories.
-This rendering process is a vital component of our experimental pipeline, enabling us to visualize and analyze the uncertainty predictions alongside the scene geometry.
+This rendering process is a vital component of our experimental pipeline, enabling us to visualize and analyze uncertainty predictions alongside scene geometry.
 The method supports producing multiple outputs, most importantly: RGB images, Depth images (visualized with a colormap), and Uncertainty maps (also colormapped, typically “inferno”).
 Each pixel’s uncertainty value is fed through a user-specified color mapping function to produce interpretive 2D heatmaps of the 3D uncertainty field.
-Figure 1: Africa Dataset BayesSDF Uncertainty Rendering (Left: RGB, Middle: Uncertainty, Right: Depth)
-Figure 2: Basket Dataset BayesSDF Uncertainty Rendering (Left: Depth, Middle: Uncertainty, Right: RGB)
+Figure 1: Africa Dataset: BayesSDF-derived view with uncertainty rendering. The left panel shows the RGB image, the middle panel depicts predictive uncertainty, and the right panel illustrates depth.
+Figure 2: Basket Dataset: BayesSDF-generated view with uncertainty.
+The left panel displays the depth map, the middle panel shows the uncertainty estimation, and the right panel presents the RGB image.
 The renderer exposes multiple camera path generation strategies:
@@
 Below are the key metrics, illustrated in our figures and tables.
-Figure 3: Africa Dataset $\Delta\text{MAE}$ Sparsification Curve
-Figure 4: Basket Dataset $\Delta\text{MAE}$ Sparsification Curve
-Figure 5: Africa Dataset Ensemble Vs. BayesSDF Method Comparison
-Figure 6: Basket Dataset Ensemble Vs. BayesSDF Method Comparison
+Figure 3: Africa Dataset: Change in mean absolute error ($\Delta$MAE) as a function of sparsification.
+Figure 4: Basket Dataset: Sparsification curve showing the change in mean absolute error ($\Delta$MAE) with increasing levels of sparsification.
+Figure 5: Africa Dataset: Comparison of the Ensemble method and BayesSDF in terms of AUSE.
+Figure 6: Basket Dataset: Comparison of the Ensemble method and BayesSDF in terms of AUSE.
   • Pixel Sparsification (x-axis in our plots): We sort the pixels in each rendered view from most uncertain to least uncertain and remove them in cumulative stages.
@@
 Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors.
-If the model is correct that “most uncertain” pixels also have high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
-  • Area Under the Sparsification Error (AUSE) with MAE: We integrate $\Delta\text{MAE}$ across the entire sparsification curve.
-A lower AUSE value indicates better alignment between “predicted uncertainty” and “true error,” meaning the model is effectively ranking pixels by their likelihood of error.
+If the model correctly identifies ”most uncertain” pixels as having high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
+  • Area Under the Sparsification Error (AUSE) with MAE: We integrate $\Delta\text{MAE}$ over the entire sparsification curve.
+A lower AUSE value indicates better alignment between “predicted uncertainty” and “true error,” indicating the model is effectively ranking pixels by their likelihood of error.
 The figures show three colored lines for each dataset/view: the purple line represents sorting by BayesSDF’s uncertainty in descending order, the blue line represents sorting by variance or alternate uncertainty measure (e.g., an ensemble-based variance), and the red line represents the difference between the two curves’ $\Delta\text{MAE}$.
 When the red line reaches a low value, it implies that our Hessian-based uncertainty is identifying the same high-error pixels that the variance-based approach finds.
-6.3 Results and Analysis
+6.3 Analysis
 As we navigate the scene via spiral or interpolated trajectories, the uncertainty maps generated by BayesSDF reveal hotspots in regions with fine geometry (e.g., object boundaries, reflective surfaces) or limited camera coverage.
@@
 BayesSDF 0.218 0.268 0.148 0.182
-Table 2: Area Under the Sparsification Error (AUSE) with Mean Absolute Error (MAE) across different datasets and methods. Lower values indicate better performance.
+Table 2: Area Under the Sparsification Error (AUSE) with Mean Absolute Error (MAE) for various datasets and reconstruction methods. Lower AUSE values indicate superior performance.
   • Lower AUSE MAE implies higher correlation between uncertainty and depth error.
@@
   1. 1.
-    Single-Pass Hessian Approximation: Rather than training multiple models (ensembles) or repeatedly sampling dropouts, BayesSDF captures uncertainty by analyzing local color-space derivatives with respect to the scene’s underlying deformation field.
+    Single-Pass Hessian Approximation: Rather than training multiple models (ensembles) or repeatedly sampling dropouts, BayesSDF captures uncertainty by analyzing local colorspace derivatives with respect to the scene’s underlying deformation field.
 This approach is both memory and compute efficient.
   2. 2.
-    Geometry-Driven Uncertainty: By tying uncertainty with signed distance function’s sensitivity, BayesSDF provides a physically interpretable metric.
-Regions where tiny deformation changes provoke large color changes correspond exactly to geometry that is poorly constrained or ambiguous.
+    Geometry-Driven Uncertainty: BBy relating uncertainty to the signed distance function’s sensitivity, BayesSDF provides a physically interpretable metric.
+Regions where small deformation changes cause large color changes correspond to geometry that is poorly constrained or ambiguous.
   3. 3.
@@
 We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage.
-These findings suggest that our Laplace-style approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
+These findings suggest that our Laplace approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
 7 Applications
@@
 BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics.
-By leveraging ensemble-based uncertainty, BayesSDF quantifies spatial ambiguities in reconstructed regions.
-These uncertainty maps enable robotic systems to detect areas of high ambiguity, which is invaluable for autonomous navigation in dynamic or cluttered environments.
+BayesSDF quantifies spatial ambiguities in reconstructed regions.
+These uncertainty maps enable robotic systems to detect areas of high ambiguity, invaluable for autonomous navigation in dynamic or cluttered environments.
 For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements.
-Physics-based simulations in fields such as engineering and medical imaging rely heavily on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution.
-BayesSDF enhances these simulations by providing uncertainty-aware geometric models.
+Physics-based simulations in fields such as engineering and medical imaging rely on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution.
+BayesSDF enhances these simulations by providing geometric models with quantified uncertainty.
 For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools.
@@
 Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs, and our method builds on this advantage to provide meaningful geometric uncertainty.
-The result is a compact, computationally efficient uncertainty estimator that aligns uncertainty directly with surface quality.
+The result is a compact, efficient uncertainty estimator that aligns uncertainty directly with surface quality.
 By utilizing models like NeuS’ volumetric SDF space, BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification.
@@
 Through rigorous experiments on synthetic and real-world datasets, BayesSDF demonstrates outstanding performance, achieving very high calibration results and reduced computational overhead.
-Looking forward, the principles of BayesSDF can be extended to broader applications, including dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations.
+Looking forward, the principles of BayesSDF can be extended to broader applications, such as dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations.
 Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks presents exciting opportunities for real-time 3D uncertainty quantification.

### v3

@@
 Abstract
-Quantifying uncertainty in neural implicit 3D representations, particularly those utilizing Signed Distance Functions (SDFs), remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies.
-Existing methods typically neglect direct geometric integration, leading to poorly calibrated uncertainty maps.
-We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and reliable uncertainty estimates are essential.
-Unlike radiance-based models such as Neural Radiance Fields (NeRF) or 3D Gaussian splatting, which lack explicit surface formulations, Signed Distance Functions (SDFs) define continuous and differentiable geometry, making them better suited for physical modeling and analysis.
-BayesSDF leverages a Laplace approximation to quantify local surface instability using Hessian-based metrics, enabling efficient, surfaceaware uncertainty estimation.
-Our method shows that uncertainty predictions correspond closely with poorly reconstructed geometry, providing actionable confidence measures for downstream use.
-Extensive evaluations on synthetic and real-world datasets demonstrate that BayesSDF outperforms existing methods in both calibration and geometric consistency, establishing a strong foundation for uncertainty-aware 3D scene reconstruction, simulation, and robotic decision-making.
+Accurate surface estimation is critical for downstream tasks in scientific simulation, and quantifying uncertainty in implicit neural 3D representations still remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies.
+However, current neural implicit surface models do not offer a principled way to quantify uncertainty, limiting their reliability in real-world applications.
+Inspired by recent probabilistic rendering approaches, we introduce BayesSDF, a novel, probabilistic framework for uncertainty estimation in neural implicit 3D representations.
+Unlike radiance-based models such as Neural Radiance Fields (NeRF) or 3D Gaussian Splatting, Signed Distance Functions (SDFs) provide continuous, differentiable surface representations, making them especially well-suited for uncertainty-aware modeling.
+BayesSDF applies a Laplace approximation over SDF weights and derives Hessian-based metrics to estimate local geometric instability.
+We empirically demonstrate that these uncertainty estimates correlate strongly with surface reconstruction error across both synthetic and real-world benchmarks.
+By enabling surface-aware uncertainty quantification, BayesSDF lays the groundwork for more robust, interpretable, and actionable 3D perception systems.
 1 Introduction
-Uncertainty quantification is a fundamental challenge in machine learning, particularly in 3D scene reconstruction, where ensuring reliability in learned representations is crucial for downstream applications such as robotics, medical imaging, and physical simulations.
-Neural implicit representations have revolutionized 3D geometry learning by encoding scenes as continuous functions rather than discrete point clouds or meshes.
-Among these, Neural Radiance Fields (NeRF) [5] enable photorealistic novel view synthesis but lack explicit geometric constraints, making them less suitable for precise surface reconstruction.
-To address this limitation, Neural Implicit Surfaces (NeuS) [7] employ Signed Distance Functions (SDFs), providing a more structured representation of 3D geometry by modeling surfaces as zero-level sets of a continuous field.
-This geometric formulation allows NeuS to achieve superior surface consistency and geometric regularity, but the challenge remains in effectively quantifying uncertainty in the learned SDF field.
-In this paper, we introduce BayesSDF, a novel probabilistic framework that enables surface-aligned uncertainty estimation in SDF-based 3D neural representations.
-Rather than relying on ensembles or pixel-level sampling, BayesSDF utilizes a Laplace approximation over a learned deformation field to compute a Hessian-based metric that captures the local sensitivity of the rendering function.
-This formulation ensures that uncertainty maps are geometrically coherent.
-By integrating the geometric advantages of SDF-enabled models with a principled probabilistic framework, BayesSDF establishes an unexplored view for uncertainty-aware 3D scene reconstruction and representation learning.
-The method enables precise identification of areas with high reconstruction variance, leading to improved surface-aware uncertainty estimates.
-BayesSDF also improves calibration and geometric consistency over previous methods.
-Instead of estimating uncertainty through pure pixel-space renderings or depth-based sampling like NeRF, BayesSDF utilizes the continuous SDF field from geometrically adaptable reconstruction models, reducing computational overhead while maintaining accuracy.
+Scientific applications such as fluid simulation, medical reconstruction, and robotic perception require not only accurate surface reconstructions but also reliable uncertainty estimates [1] to support robust decision-making (e.g.
+aleatoric vs epistemic [14]).
+Recent advances in multi-view 3D reconstruction using neural implicit representations [20] have dramatically improved surface quality, yet these methods are rarely adopted in the real world [19] due to their inability to model surface uncertainty.
+Neural implicit representations model geometry as continuous functions [20, 3, 4], enabling high-resolution reconstructions without relying on dense voxels or explicit meshes.
+While these methods, like Neural Radiance Fields [20], are effective at view synthesis, they lack surface-level structure and provide no mechanism for quantifying geometric uncertainty.
+Models like NeuS [47] (and dynamic topological models like Nerfies [30], HyperNeRF [31], CamP [park2022campnerf], etc.) address this limitation by using Signed Distance Functions (SDFs), which define surfaces as zero-level sets of continuous scalar fields.
+This provides strong geometric structure and differentiability, enabling more consistent and physically meaningful reconstructions.
+However, these models remain deterministic and offer no mechanism for estimating uncertainty over the learned surface SDF fields.
+To fill this gap, we introduce BayesSDF, a probabilistic framework for estimating surface-aligned uncertainty in neural implicit models that use SDFs.
+Our work is inspired by BayesRays [10], which introduced uncertainty modeling for volumetric radiance fields via Bayesian rendering.
+However, BayesSDF departs fundamentally from BayesRays by focusing on explicit surface geometry.
+Instead of modeling uncertainty through pixel-space or volumetric sampling, BayesSDF leverages a Laplace approximation [9] over SDF weights to quantify local surface sensitivity using second-order derivatives.
+This results in uncertainty maps that are geometrically coherent and directly tied to surface stability.
+By combining the structure of SDF-based models with a tractable probabilistic approximation, BayesSDF produces uncertainty estimates that strongly correlate with reconstruction error, offering more interpretable and actionable outputs.
+These estimates help identify ambiguous or ill-constrained geometry, which is critical for downstream tasks that require reliable spatial reasoning [50].
+Unlike uncertainty models based on pixel-space rendering or volumetric sampling, BayesSDF provides surface-aware estimates rooted in the geometry itself.
+This enables more efficient and scalable modeling without sacrificing spatial precision.
+To summarize, our key contributions are as follows:
+  • Surface-Aware Uncertainty via SDF Geometry: Prior methods often model uncertainty indirectly in pixel or depth space, which can misalign with true geometric errors.
+BayesSDF instead leverages the analytic gradients of Signed Distance Functions to estimate uncertainty directly on the surface.
+This direct integration enables uncertainty estimates that are physically meaningful and geometrically consistent.
+  • Uncertainty Estimation through Local Surface Sensitivity: BayesSDF uses a Laplace approximation and Hessian-based calculations over the neural SDF weights, capturing how local geometric perturbations affect surface stability in the geometry’s deformation field.
+By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty, avoiding the computational expense of traditional ensemble [17] or extensive sampling approaches.
+  • Scalable Deformation Modeling with Hash Encoding: To model local geometric variation, BayesSDF introduces a hierarchical multi-resolution hash-encoded deformation field.
+Compared to grid-based representations used in prior probabilistic methods like BayesRays [10], our approach supports higher spatial resolution with significantly reduced memory overhead.
+This design makes BayesSDF scalable to complex scenes while remaining compatible with frameworks like NeuS [47].
+Motivated by computational physics and scientific simulation applications with 3D environments such as modeling fluid flow through forests, where precise surface geometry and reliable uncertainty estimates are essential.
+SDF-based models, as demonstrated in DeepSDF [29], provide surface continuity and differentiability that are advantageous for physical modeling.
 The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling.
-This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
-2 Contributions
+Our framework explicitly supports such use cases by relating uncertainty to local surface instability.
+2 Related Work
+Uncertainty estimation is essential in 3D learning systems, especially for safety-critical and scientific applications, where downstream tasks depend on the reliability of reconstructed geometry.
+Traditional neural implicit methods such as Neural Radiance Fields (NeRF) [20, 19] and Signed Distance Function (SDF)-based models like NeuS [47] have achieved high-fidelity reconstruction [59, 44] and view synthesis, but largely ignore uncertainty modeling.
+This omission limits their usefulness in tasks requiring interpretable or risk-aware decision-making.
+In this section, we contextualize BayesSDF, within three major threads of prior work:
   1. 1.
-    Direct Geometric Integration: Unlike existing techniques that indirectly approximate uncertainty through pixel-space or depth-based approaches, BayesSDF explicitly integrates the geometric properties inherent in Signed Distance Functions.
-This direct integration significantly improves the geometric consistency and calibration of uncertainty predictions.
+    Why SDF-based models [29, 47] offer superior structure for geometric uncertainty estimation compared to NeRF-style density fields. [20, 19, 7]
   2. 2.
-    Efficient Hessian-based Uncertainty Estimation: BayesSDF employs a Hessian-based uncertainty calculation method that efficiently captures local sensitivity in the geometry deformation field.
-By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty, avoiding the computational expense of traditional ensemble or extensive sampling approaches.
+    How uncertainty has been previously modeled in neural 3D systems, including ensemble methods and Bayesian inference.
   3. 3.
-    Scalable and Computationally Efficient Framework: Leveraging a hierarchical hash-based deformation field, BayesSDF maintains high accuracy while reducing computational and memory overhead.
-The approach integrates seamlessly with existing neural implicit surface frameworks such as NeuS, ensuring a compact yet powerful uncertainty estimation model.
-  4. 4.
-    Comprehensive Evaluation and Validation: Extensive empirical validation on synthetic and real-world datasets demonstrates BayesSDF’s superiority over existing uncertainty quantification methods.
-Metrics such as the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially in complex real-world scenarios.
-  5. 5.
-    Scientific Motivation for SDF-Based Uncertainty: We motivate our approach by targeting applications in scientific simulation and computational physics, where accurate geometry and interpretable uncertainty are critical.
-SDF-based models, as demonstrated in DeepSDF [6], provide surface continuity and differentiability that are advantageous for physical modeling.
-Our framework explicitly supports such use cases by relating uncertainty to local surface instability.
-These contributions establish BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications for scene reconstruction, physics-based simulations, and autonomous robotic navigation.
-3 Related Work
-Uncertainty quantification in neural implicit representations is a rapidly evolving field.
-Traditional deep learning models often produce deterministic outputs without explicit confidence measures, making them unreliable in critical scenarios.
-In the context of 3D learning, uncertainty quantification plays a crucial role in assessing reconstruction accuracy, identifying ambiguous regions, and guiding active learning.
-Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, among others.
-We then highlight how BayesSDF significantly advances the field by integrating uncertainty directly into SDF-based neural implicit models, enabling computationally efficient and geometrically consistent uncertainty estimation.
+    Why BayesSDF provides a principled and non-trivial improvement over prior Bayesian extensions like BayesRays [10], particularly in aligning uncertainty with surface structure.
 2.1 Uncertainty Estimation in 3D Reconstruction Methods
-Uncertainty estimation plays a crucial role in 3D reconstruction, where accurate surface modeling is essential for downstream applications.
-Various methods exist for reconstructing 3D surfaces, and each presents unique challenges for uncertainty quantification.
+3D reconstruction methods vary widely in how they represent geometry, which directly impacts how uncertainty can be quantified.
+NeRF-style models [20, 3, 4] rely on volume densities sampled along camera rays.
+These are well-suited for view synthesis but decouple uncertainty from physical surfaces.
+In contrast, SDF-based models [29, 39, 36, abou2022freenerf, 40, 12] define surfaces as the zero level-set of continuous scalar fields, offering geometric interpretability and a natural differentiable structure [28, 33, 54], which are features critical for scientific computing tasks involving simulations, collisions, or boundary conditions.
+BayesSDF builds on this structure by estimating uncertainty not in ray-space but in the SDF’s parameter space.
+This enables surface-aware uncertainty aligned with real geometric features rather than visual artifacts.
 2.1.1 Neural Representations and SDF-Based Methods
-Neural implicit representations, such as NeRF (Neural Radiance Fields) and NeuS (Neural Implicit Surfaces), have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction.
+Neural implicit representations, such as NeRF [20] and NeuS [47], have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction.
 However, their uncertainty estimation techniques differ significantly.
-  • NeRF-based methods estimate uncertainty per-ray by analyzing the variance in color and density outputs across multiple samples along a viewing ray.
-These approaches are effective for rendering applications but fail to capture geometric uncertainty directly.
-  • NeuS-based methods instead represent surfaces using a SDF, which provides a continuous and differentiable description of 3D geometry.
-However, existing NeRF uncertainty estimation techniques do not translate well to NeuS because they rely on ray-space density variations rather than surface-level geometric uncertainty.
-3.2 Uncertainty Estimation in Ensemble-Based Learning
-Ensemble-based uncertainty estimation has emerged as a powerful technique in deep learning, providing a practical means of approximating epistemic uncertainty.
-In the context of 3D learning and reconstruction methods, ensemble methods are particularly advantageous as they offer robust uncertainty quantification without requiring complex probabilistic modeling or extensive sampling.
+  • NeRF-based methods [20, 37, 38] estimate the uncertainty per ray by analyzing the variance in color and density outputs across multiple samples along a viewing ray.
+These approaches are effective for rendering applications, but they do not capture geometric uncertainty directly [2].
+  • NeuS-based methods [47] instead represent surfaces using an SDF, which provides a continuous and differentiable description of 3D geometry.
+[41, 13] However, existing NeRF [20] uncertainty estimation techniques do not translate well to NeuS [47] because they rely on ray-space density variations rather than surface-level geometric uncertainty.
+2.2 Ensemble-Based Methods and Their Limitations
+Ensemble-based methods [17, 1, 14] are a standard approach to approximate epistemic uncertainty by training and sampling multiple networks.
+In the context of reconstruction methods, ensemble methods are particularly advantageous as they offer robust uncertainty quantification without requiring complex probabilistic modeling or extensive sampling.
 While ensemble methods are widely used in 2D vision tasks (e.g.
-image classification and segmentation), applying them effectively in 3D neural representations presents unique challenges:
-  • Computational Cost: Training multiple neural implicit models is memory- and computeintensive, particularly for high-resolution 3D reconstructions.
-Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues.
-  • Surface Consistency: Unlike 2D image-based tasks, uncertainty in 3D learning must be estimated consistently across a continuous surface rather than at discrete pixel locations
+image classification and segmentation), they introduce several limitations in the 3D domain:
+  • Scalability: Training multiple neural implicit models is memory and compute intensive, particularly for high-resolution 3D reconstructions.
+[23, 8, 55, 45] Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues [48].
+  • Surface Consistency: Unlike 2D image-based tasks [21], uncertainty in 3D learning must be estimated consistently across a continuous surface rather than at discrete pixel locations. [61, 58]
   • Ray vs.
-Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like models operate strictly along rays.
+Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like [20] models operate strictly along rays.
 This results in uncertainty that does not align well with geometric structures, making it less interpretable.
-While leveraging the strengths of ensemble-based learning and addressing key limitations in 3D implicit representations, BayesSDF establishes a new paradigm for uncertainty quantification in neural implicit fields.
+These challenges motivate more principled Bayesian approximations within neural 3D representations leveraging the strengths of ensemble-based learning, particularly those offering geometric continuity like SDFs [53].
 2.3 Bayesian Inference for Ray-Based Sampling
-BayesRays [2] is a recent approach for uncertainty quantification in neural radiance fields.
-It modifies NeRF by introducing a Bayesian framework that estimates uncertainty in the color and density outputs of the model.
-While BayesRays is effective for quantifying NeRF uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty.
-Additionally, BayesRays relies on ray-space sparsification techniques, which can lead to misalignment between estimated uncertainty and actual geometric uncertainty.
-Unlike BayesRays, BayesSDF integrates reconstruction methods that inherently reveal volumetric qualities, ensuring direct alignment between uncertainty estimates and the scene geometry.
-This approach places less emphasis on rendering estimations, making it inherently more interpretable for geometric reconstruction tasks by exploiting surface variability rather than depth-based uncertainty.
+BayesRays [10] introduced uncertainty estimation in NeRF [20] by modeling pixel-space color and density distributions using a Bayesian ensemble and ray sparsification.
+While BayesRays [10] is effective for capturing radiance uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty [37, 32, 6, 60, 34, 11].
+It relies on ray-space sparsification techniques [30, 31, 25, 46], which can lead to misalignment between estimated uncertainty and actual geometric uncertainty [43].
+Its volumetric density representation lacks an explicit notion of surface, making it harder to isolate geometric ambiguity.
+BayesSDF addresses these issues by modeling the sensitivity of the surface geometry itself.
+Rather than relying on variation in rendered pixel values, BayesSDF quantifies how local perturbations in the SDF field affect surface stability using a Laplace approximation [9, 39, 36, 26, 52, 35].
+This results in uncertainty maps that are tightly coupled to the surface [29, 24], making them more interpretable and physically meaningful.
+Notably, directly applying BayesRays [10] to SDF-based models like NeuS [47] is non-trivial.
+BayesRays [10] relies on sampling distributions along rays [29, 27], which do not generalize to SDF’s zero-crossing surface logic.
+In contrast, BayesSDF incorporates the analytic structure of SDFs via second-order derivatives [9, 39, 24], enabling uncertainty estimates that are both tractable and geometrically consistent.
 Method Representation Uncertainty Estimation Computational Cost Geometric Consistency
-NeRF Volume Density None High Low
-NeuS SDF None High High
-BayesRays Volume Density Ensemble Variance Low Moderate
-BayesSDF Volume Density Ensemble Variance Low Very High
-Table 1: Comparison of existing 3D reconstruction methods, evaluating their representation, uncertainty estimation, computational cost, and geometric consistency.
+NeRF [20] Volume Density None High Low
+NeuS [47] SDF None High High
+BayesRays [10] Volume Density Ray-Space Variance Moderate Moderate
+BayesSDF SDF Hessian-Based Laplace Approx. Low Very High
+Table 1: Comparison of existing 3D reconstruction methods based on geometry representation, uncertainty modeling, computational cost, and surface-level interpretability.
 3 Background
-Quantifying uncertainty in 3D implicit representations is essential for improving the reliability and interpretability of learned geometric models.
-This section reviews the foundational components upon which BayesSDF builds: Signed Distance Functions (SDFs), Neural Radiance Fields (NeRF), Neural Implicit Surfaces (NeuS), ensemble-based uncertainty methods, and Bayesian estimation frameworks such as BayesRays.
-4.1 Signed Distance Functions (SDFs)
+Quantifying uncertainty in 3D implicit representations is essential to improve the reliability and interpretability of learned geometric models.
+This section reviews the foundational components upon which BayesSDF builds upon.
+3.1 Neural Implicit Surfaces (NeuS)
 Signed Distance Functions (SDFs) define geometry as a continuous field that measures the signed distance of a point in space to the closest surface boundary.
 Formally, an SDF is defined as $f:\mathbb{R}^{3}\to\mathbb{R}$, where:
-$f(\mathbf{x})=\begin{cases}\phantom{-}d,&\text{if }\mathbf{x}\text{ is outsidethe surface}\\ \phantom{-}0,&\text{if }\mathbf{x}\text{ lies on the surface}\\ -d,&\text{if }\mathbf{x}\text{ is inside the surface}\end{cases}$
-This representation allows for accurate computation of surface normals and curvature, and forms the basis of many modern 3D reconstruction methods [6].
-Despite their geometric advantages, SDFs have not been fully explored for probabilistic modeling or uncertainty estimation in learned 3D representations.
-4.2 Neural Radiance Fields (NeRF)
-NeRF models learn a volumetric function that maps 3D positions and viewing directions to emitted color and volume density [5].
-The function $F_{\theta}:(\mathbf{x},\mathbf{d})\rightarrow(\sigma,\mathbf{c})$ enables photorealistic novel view synthesis by rendering images through volume integration:
 […]
-[…]
-4.3 Neural Implicit Surfaces (NeuS)
-NeuS addresses the geometric limitations of NeRF by replacing volumetric density fields with SDFs that model surfaces as zero-level sets [7].
+This representation allows for accurate computation of surface normals and curvature, and forms the basis of many modern 3D reconstruction methods [29, 5, 12, 33, 28, 54, 26].
+Despite their geometric advantages, SDFs have not been fully explored for probabilistic modeling or uncertainty estimation in learned 3D representations [18, 11, 53].
+NeuS [47] addresses the geometric limitations of NeRF [20] by replacing volumetric density fields with SDFs that model surfaces as zero-level sets.
 The rendering is adapted by transforming SDF values into opacity via a sigmoid function:
@@
 $C(\mathbf{r})=\sum_{i}T_{i}\alpha_{i}\mathbf{c}_{i},\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$
-NeuS enables precise geometric modeling, but lacks inherent mechanisms for uncertainty quantification.
+NeuS [47] enables precise geometric modeling, but lacks inherent mechanisms for uncertainty quantification.
 3.2 Uncertainty Estimation via Ensembles
@@
 $\bar{f}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}f_{i}(\mathbf{x}),\quad\text{Var}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}(f_{i}(\mathbf{x})-\bar{f}(\mathbf{x}))^{2}$
-This variance captures model disagreement, which is indicative of uncertainty [4].
-While ensemble methods are effective across 2D and 3D vision tasks [1], they are resource-intensive and can be misaligned with underlying geometry in volumetric or ray-based models.
-4.5 Bayesian Uncertainty in NeRF: BayesRays
-BayesRays proposes a Bayesian approach for NeRF by learning a density-aware variance model and propagating it along rays [2]. The estimated uncertainty is given by:
-$U(\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}(\mathbf{x})\cdot T_{k}(\mathbf{x})$
-Color and density are modeled as Gaussian random variables, and uncertainty in the rendered image is computed by propagating these distributions through the rendering equation:
-$\text{Var}[\hat{C}(\mathbf{r})]=\sum_{i=1}^{N}T_{i}^{2}\left(1-\exp(-\sigma_{i}\delta_{i})\right)^{2}\text{Var}[\mathbf{c}_{i}]$
-BayesRays achieves efficient uncertainty estimation without ensembles, but its output lacks surface-level geometric consistency.
-Integrating such uncertainty with geometric fields remains a core challenge in the field [3].
+This variance captures model disagreement, which is indicative of uncertainty [17, 14, 2, 57].
+While ensemble methods are effective across 2D and 3D vision tasks [1], they are resource-intensive and can be misaligned with underlying geometry in volumetric or ray-based models [61, 56].
 4 Method
-In this section, we detail our proposed BayesSDF method, which leverages a NeuS-like scene representation and adopts an uncertainty estimation framework through a deformation field and a Hessian-based metric.
-We begin by describing the theoretical underpinnings and motivation, then present how we construct and query the deformation field, calculate the Hessian, and finally approximate local uncertainty.
+In this section, we detail our proposed BayesSDF method, which leverages a NeuS [47] scene representation and adopts an uncertainty estimation framework through a Hessian-based deformation field.
+We begin by describing the theoretical underpinnings, then present how we construct and query the deformation field, calculate the Hessian, and finally approximate local uncertainty.
 4.1 Theory
-Our goal is to model an associated uncertainty measure, derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters.
-Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in Red Green Blue (RGB)) rendered by the neural field.
-Rather than relying purely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field.
+Our goal is to model an associated uncertainty measure derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters.
+Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in the scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in red-green-blue (RGB)) rendered by the neural field.
+Rather than relying solely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field [32, 30].
 Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$.
-To estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$.
-Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction.
+To estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$ [14, 1].
+Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction [9, 35].
 4.2 Distortion Fields
-The distortion field $f_{\mathrm{hash}}(\mathbf{x})$ captures spatial distortions in the SDF field, serving as a basis for uncertainty estimation.
-To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding that functions as a voxel-like grid implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, linear interpolation between grid corners, and 3 output channels per queried grid vertex.
-Effectively, for a given point $\mathbf{x}$, the hash encoding looks up an 8-corner neighborhood around the voxel containing $\mathbf{x}$, interpolates among these corners (each storing a 3D deformation vector), and outputs a single 3D vector $\mathbf{d}(\mathbf{x})$.
-This approach is parameter-efficient while remaining flexible.
-We denote $\mathbf{aabb}\in\mathbb{R}^{6}$ as the scene bounding box, which is used to normalize $\mathbf{x}$ to a unit cube or appropriate domain before querying the hash grid.
-Specifically, we generate a ray bundle from the dataset’s rays, containing origins, directions, and near/far bounds.
-We then use Neus’ hierarchical or “proposal” sampler to refine sample locations along each ray, focusing more samples near scene geometry (i.e., near the SDF zero level set).
-Let $\{\mathbf{x}_{i}\}$ be the set of sampled 3D points on a ray.
-When each ray sample $\mathbf{x}\in\mathbb{R}^{3}$ is generated, we transform $\mathbf{x}$ into normalized coordinates $\tilde{\mathbf{x}}=\mathrm{normalize}(\mathbf{x},\mathbf{aabb})$.
-Then, $\mathbf{d}(\mathbf{x})\;=\;\sum_{c=1}^{8}\alpha_{c}(\tilde{\mathbf{x}})\,\mathbf{d}_{c}\quad\text{where}\quad\bigl{\{}\mathbf{d}c\bigr{\}}{c=1}^{8}\;$ are are the corner deformation vectors that are found from a grid cell index in each dimension (one for each binary combination of $\lfloor\cdot\rfloor or\lceil\cdot\rceil$ in (x,y,z)) and $\alpha_{c}(\tilde{\mathbf{x}})$ are the trilinear interpolation coefficients $(\alpha_{c}(\mathbf{x}i)\;\in\;[0,1],\quad c=1,\ldots,8,)$ such that $\sum{c=1}^{8}\alpha_{c}(\mathbf{x}_{i})=1$.
-These coefficients are used to blend the deformation vectors from the 8 corners.
-This procedure is vital not only for generating $\mathbf{d}(\mathbf{x}_{i})$, but also for accumulating partial derivatives.
-Because each point’s deformation influences eight corner entries, we track which corners are involved and in what proportion.
+[…]
 4.3 Calculating the Hessian
@@
 Concretely, for each corner’s index k in the hash grid, the squared sum of partial derivatives is computed as follows:
-$\Delta(\mathbf{x})\;=\;\sum_{d=1}^{3}\Bigl{(}\frac{\partial r}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial g}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial b}{\partial\mathbf{d}_{d}}^{2}\Bigr{)},$
+$\Delta(\mathbf{x})\;=\;\sum_{d=1}^{3}\left(\left(\frac{\partial r}{\partial\mathbf{d}_{d}}\right)^{2}\cdot\left(\frac{\partial g}{\partial\mathbf{d}_{d}}\right)^{2}\cdot\left(\frac{\partial b}{\partial\mathbf{d}_{d}}\right)^{2}\right)$
 where $\mathbf{d}_{d}\in\{\mathbf{d}_{1},\mathbf{d}_{2},\mathbf{d}_{3}\}$ represents each dimension of the deformation vector.
@@
 We interpret these aggregated squared gradients in ‘hessian‘ as proxies for uncertainty.
-High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color, suggesting the reconstruction at that location is sensitive—and potentially uncertain.
+High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color [17, 24, 18], suggesting the reconstruction at that location is sensitive and potentially uncertain [11, 10].
 Conversely, near-zero partial derivatives indicate that changes in local deformation parameters barely affect the color, suggesting more stable or confident estimates.
@@
 Below are the key metrics, illustrated in our figures and tables.
-Figure 3: Africa Dataset: Change in mean absolute error ($\Delta$MAE) as a function of sparsification.
-Figure 4: Basket Dataset: Sparsification curve showing the change in mean absolute error ($\Delta$MAE) with increasing levels of sparsification.
+[…]
+[…]
 Figure 5: Africa Dataset: Comparison of the Ensemble method and BayesSDF in terms of AUSE.
@@
     where $\hat{d}(u,v)$ is the predicted depth, $d^{(}u,v)$ is ground-truth depth, and $\Omega(p)$ is the set of remaining pixels after removing a fraction p of the domain based on uncertainty sorting.
-Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors.
+Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors [14, 1].
 If the model correctly identifies ”most uncertain” pixels as having high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
@@
   • Stable Surfaces (with consistent multi-view coverage) often exhibit low uncertainty, reinforcing that the Hessian-based gradients reflect reconstruction confidence.
-[…]
+These visualizations serve as a powerful diagnostic for identifying challenging areas and potential modeling failures.
+In the sparsification curves (Figure $\Delta\text{MAE}\text{ vs.
+}\text{Pixel Sparsification}$), both BayesSDF’s purple line and the variance-based blue line converge to similar cumulative MAE values as the sparsification approaches the least uncertain pixels.
+This convergence suggests that BayesSDF’s uncertainty ranking effectively captures error-prone pixels, similar to variance-based or ensemble-based uncertainty techniques [17, 24, 57, 18].
+We present a table that summarizes the Area Under the Sparsification Error curve (with Mean Absolute Error) across multiple scenes:
 Method Basket Africa Statue Torch
@@
   • BayesSDF matches or slightly exceeds ensemble methods on several scenes (notably, “Statue”), while on “Africa” it shows a modestly higher AUSE.
-  • Despite not always having the absolute lowest metric, BayesSDF’s performance remains highly competitive while avoiding the large computational expense of ensemble training.
+  • Despite not always having the absolute lowest metric, BayesSDF’s performance remains highly competitive while avoiding the large computational expense of ensemble training [11, 10].
 Through BayesSDF, we are able to achieve:
@@
   3. 3.
-    Compact Integration into NeuS: BayesSDF only requires storing a Hessian-based statistic in the resolution-limited deformation grid, leading to a small parameter footprint while seamlessly integrating into the existing NeuS rendering pipeline.
-To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets.
+    Compact Integration into NeuS: BayesSDF only requires storing a Hessian-based statistic in the resolution-limited deformation grid, leading to a small parameter footprint while seamlessly integrating into the existing NeuS [47] rendering pipeline.
+To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets [9, 35, 48].
 We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage.
-These findings suggest that our Laplace approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
+These findings suggest that our Laplace approximation effectively identifies instability in the underlying SDF geometry [9, 35, 29, 54], making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
 6 Applications
@@
 Below, we discuss the core applications of BayesSDF, emphasizing its technical strengths and relevance across various fields.
-BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics.
+BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics [20, 23, 26, 47].
 BayesSDF quantifies spatial ambiguities in reconstructed regions.
-These uncertainty maps enable robotic systems to detect areas of high ambiguity, invaluable for autonomous navigation in dynamic or cluttered environments.
-For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements.
-Physics-based simulations in fields such as engineering and medical imaging rely on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution.
-BayesSDF enhances these simulations by providing geometric models with quantified uncertainty.
-For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools.
-BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology.
-For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels, allowing researchers to distinguish between highly reliable and uncertain reconstructions.
+These uncertainty maps enable robotic systems to detect areas of high ambiguity, invaluable for autonomous navigation in dynamic or cluttered environments [53, 58, 41, 25].
+For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements [22, 50, 39].
+Physics-based simulations in fields such as engineering and medical imaging rely on accurate 3D models [1, 14, 2] to simulate phenomena such as fluid dynamics or stress distribution [36, 42].
+BayesSDF enhances these simulations by providing geometric models with quantified uncertainty [10, 11, 18].
+For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools [51, 17, 24].
+BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology [33, 59, 60].
+For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels [13, 29, 54], allowing researchers to distinguish between highly reliable and uncertain reconstructions.
 7 Conclusion
-We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models.
-Motivated by scientific and physical simulation applications where accurate surface geometry is paramount, BayesSDF integrates a Laplace approximation technique over signed distance fields to quantify local reconstruction instability.
-Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs, and our method builds on this advantage to provide meaningful geometric uncertainty.
-The result is a compact, efficient uncertainty estimator that aligns uncertainty directly with surface quality.
-By utilizing models like NeuS’ volumetric SDF space, BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification.
-The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing, a computationally efficient Hessian-based curvature estimation, and ensemble-based epistemic uncertainty metrics.
-Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry and the high computational costs associated with traditional sampling-based approaches.
-Through rigorous experiments on synthetic and real-world datasets, BayesSDF demonstrates outstanding performance, achieving very high calibration results and reduced computational overhead.
-Looking forward, the principles of BayesSDF can be extended to broader applications, such as dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations.
-Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks presents exciting opportunities for real-time 3D uncertainty quantification.
+We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models [29, 26, 47].
+Motivated by scientific and physical simulation applications where accurate surface geometry is paramount [41, 19], BayesSDF integrates a Laplace approximation technique [35, 9, 24] over signed distance fields to quantify local reconstruction instability.
+Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs [20, 54, 28], and our method builds on this advantage to provide meaningful geometric uncertainty [14, 2, 1].
+The result is a compact, efficient uncertainty estimator that aligns uncertainty directly with surface quality [18, 11].
+By utilizing models like NeuS’ volumetric SDF space [47, 16], BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification.
+The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing [23], a computationally efficient Hessian-based curvature estimation [43], and ensemble-based epistemic uncertainty metrics [17, 57].
+Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry [27, 56] and the high computational costs associated with traditional sampling-based approaches [45, 8].
+Through rigorous experiments on synthetic and real-world datasets [4, 33, 49], BayesSDF demonstrates outstanding performance, achieving very high calibration results [39, 51] and reduced computational overhead.
+Looking forward, the principles of BayesSDF can be extended to broader applications, such as dynamic scene understanding [30, 32, 25, 46] and hybrid models that combine SDFs with density-based neural representations [yariv2021volume, 7].
+Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks [55, 15] presents exciting opportunities for real-time 3D uncertainty quantification.
 8 Acknowledgments